OpenStudy (anonymous):

let f(x)=(1/2)x^2. The fraction (f(a+h)-f(a))/h is?

OpenStudy (anonymous):

\[ \frac{\frac{1}{2} (a+h)^2-\frac{a^2}{2}}{h} \] Expand and simplify

OpenStudy (anonymous):

\[ \frac{\frac{a^2}{2}+a h+\frac{h^2}{2}-\frac{a^2}{2 }}{h} \]

OpenStudy (anonymous):

\[ \frac{a h+\frac{h^2}{2}}{h}= a +\frac h2 \]

OpenStudy (anonymous):

Thanks for the help